OpenStudy (anonymous):

solve for x: 12=-4/5x

OpenStudy (anonymous):

I forgot how to do this, someone please help out!

OpenStudy (anonymous):

EXAMPLE NOT YOUR MATH PROBLEM ISOLATE X VALUE 2 4 = --- 8X 2 8 [ 4 = --- ] MULTIPLY BY 8 8X 2 32 = --- X 32 = 2 / X 16 = X ( i think this example is correct)

OpenStudy (anonymous):

i think the 4 & 5 have to stay together.. -4/5 is a fraction

OpenStudy (whpalmer4):

No, they don't have to stay together. \[12 = -\frac{4}{5}x\]Multiply each side by 5, what do you get? \[5*12 = -\frac{4}{5}*5*x\]\[60 = -4x\]Now divide each side by -4 \[\frac{60}{-4} = \frac{-4}{-4}x\]\[-15 = x\] You can add, subtract, multiply, divide, take square roots, etc. so long as you do the same thing to each side of the equation.

OpenStudy (dan815):

palmer welcome to the green club

OpenStudy (dan815):

glad u could join us

OpenStudy (dan815):

ive been in the greeen club since day 1, in my heart

OpenStudy (whpalmer4):

Hey, wow! That was the question that did it! :-)

OpenStudy (dan815):

:)

OpenStudy (jhannybean):

hahaa!